given f(x)= 5x + 7 and g(x) = 5x-1 find (fg)(x)
Found 2 solutions by Fombitz, jessica43:
Answer by Fombitz(32388)
Answer by jessica43(140)

I'm assuming that (fg)(x) means composition of functions, which is usually written as (f o g)(x) or f(g(x)).
This is read as "f-compose-g of x", and means "plug x into g, evaluate, and then plug the result into f".
So if you plug in x into g(x), you get 5x-1.
Now plug that into f(x):
f(x) = 5x + 7
f(5x-1) = 5(5x-1) + 7
f(5x-1) = 25x - 5 + 7
f(5x-1) = 25x + 2
So f(g(x)) = 25x + 2